# Automatic Curriculum Design for Zero-Shot Human-AI Coordination

### v1

@@ Abstract
Zero-shot human-AI coordination is the training of an ego-agent to coordinate with humans without using human data. Most studies on zero-shot human-AI coordination have focused on enhancing the ego-agent’s coordination ability in a given environment without considering the issue of generalization to unseen environments. Real-world applications of zero-shot human-AI coordination should consider unpredictable environmental changes and the varying coordination ability of co-players depending on the environment. Previously, the multi-agent UED (Unsupervised Environment Design) approach has investigated these challenges by jointly considering environmental changes and co-player policy in competitive two-player AI-AI scenarios. In this paper, our study extends the multi-agent UED approach to a zero-shot human-AI coordination. We propose a utility function and co-player sampling for a zero-shot human-AI coordination setting that helps train the ego-agent to coordinate with humans more effectively than the previous multi-agent UED approach. The zero-shot human-AI coordination performance was evaluated in the Overcooked-AI environment, using human proxy agents and real humans. Our method outperforms other baseline models and achieves a high human-AI coordination performance in unseen environments.
@@ 1 Introduction
Deep reinforcement learning has demonstrated success across diverse domains such as gaming Silver et al. (2017); Berner et al. (2019); Vinyals et al. (2019), autonomous vehicles Kiran et al. (2021), and robotic controls Fang et al. (2019). Research on deep reinforcement learning has expanded to multi-agent reinforcement learning in which multiple agents cooperate or compete to achieve specific objectives Zhang et al. (2021); Oroojlooy and Hajinezhad (2023). Most studies in this field have focused on developing powerful AI beyond human performance through AI-AI interactions. Recently, interest has shifted to problem-solving through human-AI interaction.
@@
Some human-AI interaction studies focus on training the agents that coordinate with humans in common tasks and problem-solving. These studies typically require a large amount of human data for training but it is hard to get enough human data in the real world. To overcome this limitation, the zero-shot human-AI coordination study Strouse et al. (2021) suggests an approach that trains the ego-agent to coordinate with humans without human data. This approach mainly uses population-based training (PBT) Strouse et al. (2021); Lupu et al. (2021); Zhao et al. (2023); Lou et al. (2023). During training, the ego-agent coordinates with various co-player agents within the population pool, instead of a specific co-player agent. This prevents over-fitting to a specific coordinate strategy and helps generalize across diverse human coordinate strategies. However, these approaches only focus on unseen co-players and do not consider generalization to unseen environments. Real-world applications of human-AI coordination face various unpredictable variables and situations. Therefore, it is essential to train the ego-agent to be robust against both diverse co-players and environments.
@@
The Unsupervised environment design (UED) Dennis et al. (2020) is a method that selectively curates the environment to the agent by using a curriculum based on a given utility function. It curates increasingly difficult environments based on the current agent’s policy. This approach allows the agent to develop the ability to generalize across all possible environments by facing diverse and challenging scenarios. These methods Jiang et al. (2021a, b); Parker-Holder et al. (2022) mainly use a regret as a utility function. Recently, UED methods have been extended to competitive multi-agent settings Samvelyan et al. (2023). In a multi-agent setting, the co-player policy varies depending on the environment, so the multi-agent UED curates joint environment/co-player pair to the agent. This approach has successfully trained the robust agent across diverse environments and co-players. By extending this method to the zero-shot human-AI coordination setting, we could be able to train a robust agent that can coordinate with real humans across diverse environments.
@@
Figure 1: Overview of the proposed method. The ego-agent is trained using a population of co-players whose individual environment buffers maintain a fixed size of K, containing the environments with the lowest coordination scores. Our approach samples the co-player based on their coordination scores within the population. Depending on the replay decision, the environment is either sampled from unseen environment buffer or from the co-player’s buffer guided by a replay distribution that prioritizes the sampling of environments based on coordination scores. After sampling, our method calculates a return from the episode trajectory within the sampled environment which is utilized to compute the coordination score. This coordination score is then updated within the co-player’s buffer.
However, there are main problems with extending multi-agent UED methods to zero-shot human-AI coordination settings. The regret-based utility function and co-player sampling used in previous multi-agent UED approaches are unsuitable for these settings. These methods are designed for zero-sum games and focus on minimizing the agent’s regret to find a minimax strategy. In coordination settings, agents need to collaborate to achieve shared goals rather than minimize regret. Thus, the regret-based utility function and co-player sampling conflict with the objectives of coordination settings.
To address these problem, we introduce the automatic curriculum design for zero-shot human-AI coordination. We adopt a return-based utility function and co-player sampling for zero-shot human-AI coordination. The return measures the coordination ability of the ego-agent with the co-players. It guides the agent to get better coordination strategies even in challenging scenarios and co-players.
@@
The contributions of our study are the following: (1) We propose the multi-agent UED framework for zero-shot human-AI coordination by using a return-based utility function and co-player sampling. (2) We show that our method demonstrates higher coordination performance with human proxy agents and real human partners compared to other baselines. Figure 1 shows an overview of our proposed method.
@@ 2.1 Unsupervised Environment Design Strategies
The Unsupervised Environment Design (UED) is a method in which the teacher curates a curriculum to enable the student agent to generalize across all possible levels of the environment based on a utility function. The simplest version of UED teacher uses the utility function as a constant $C$ for uniformly random sampling the level of environment $U_{t}(\pi,\theta)=C$. Here, $t$ is the teacher, $\pi$ is the student’s policy, $\theta$ is the environment level parameter Jakobi (1997). Most UED studies use regret as a teacher’s utility function, defined as the difference between the expected return of the current policy and the optimal student policy. Teacher aims to maximize this regret Dennis et al. (2020); Gur et al. (2021), which can be defined as : $U_{t}(\pi,\theta)=\max_{\pi^{*}\in\Pi}\{\textsc{Regret}^{\theta}(\pi,\pi^{*})\}=\max_{\pi^{*}\in\Pi}\{V_{\theta}(\pi^{*})-V_{\theta}(\pi)\}$, where $\pi^{*}$ is the optimal policy on $\theta$, $\Pi$ is the set of the student policies. This regret-based utility function encourages the teacher to curate challenging environments for the student agent. Through the training process, the student’s policy will converge to the minimax regret policy by minimizing the regret in the teacher-curated environment Dennis et al. (2020): $\pi\in\arg\min_{\pi\in\Pi}\{\max_{\theta,\pi^{*}\in\Theta,\Pi}\{\textsc{Regret}^{\theta}(\pi,\pi^{*})\}\}$ where $\Theta$ is the set of environment level parameter. Since the optimal policy $\pi^{*}$ cannot be accessed in practice, regret must be approximate.
There are several UED studies on the use of regret. Prioritized Level Replay (PLR) Jiang et al. (2021b) and Robust PLR Jiang et al. (2021a) leverage a regret-based utility function to train the agent by selectively sampling environments with high learning potential. ACCEL Parker-Holder et al. (2022) continuously edits the environment that pushes the boundaries of the student agent’s capabilities by combining an evolutionary approach with Robust PLR. MAESTRO Samvelyan et al. (2023) extends single-agent UED to multi-agent UED in a competitive setting by curating the environment/co-player pairs while considering the environment-dependent co-player policy. In this paper, our method extends the competitive multi-agent UED to a zero-shot human-AI coordination setting using an appropriate utility function.
@@ 2.2 Zero-Shot Human-AI Coordination
In multi-agent cooperation tasks, achieving generalization to partner agents is an important problem. Zero-shot coordination Hu et al. (2020) and ad hoc team cooperation Stone et al. (2010) studies aim to address this problem. These approaches commonly rely on the self-play method which is an agent learns through collaboration with its clone. However, self-play methods suffer from limitations. Because they assume the partner agent to be optimal or similar to itself during the training process, its performance is poorly coordinated with unseen partners such as humans or other agents Carroll et al. (2019).
To address these issues, Devlin et al. (2011) increase the partner diversity and encourage self-play agents to learn the different strategies. One of these approaches is Population-Based Training (PBT), leveraging the diversity within a population of partner agents. In each iteration, agents are paired with different partners from the population, enabling them to learn various strategies simultaneously. Fictitious Co-Play (FCP) Strouse et al. (2021) is a well-known PBT-based approach. It uses the population of past checkpoint self-play agents as a partner to train the ego-agent. Trajectory Diversity (TrajeDi) Lupu et al. (2021) enhanced performance in cooperative tasks by generating diverse policies within the population. It used the Jensen-Shannon divergence to measure the diversity between different policies. Additionally, MEP Zhao et al. (2023) was proposed to mitigate potential issues arising from shifts in the distribution of self-play agents during interactions with unseen partners. In this study, we construct a population of co-player agents trained in various environments. Through this approach, we aim to address the poor coordination of human partners with unseen environments.
@@ 3.1 Overview
In this section, we describe our proposed method, Automatic Curriculum Design for Zero-Shot Human-AI Coordination, which is an auto curriculum that trains the ego-agent using return-based utility functions. Our method focuses on training the ego-agent to coordinate well with human partners in diverse environments, and it is based on the multi-agent UED approach Samvelyan et al. (2023). It curates environment/co-player pairs with learning potential in Human-AI coordination to the ego-agent.
@@
2:     for $N$ episodes do
3:        $\pi^{\prime}\sim\mathfrak{B}$triangleright Sample co-player via Eq.2
@@
6:           $\theta\sim\bm{\Lambda}(\pi^{\prime})$triangleright Sample a replay env with dist.1
@@
10:           $\theta\sim\Lambda_{unseen/train}$triangleright Sample unseen random env
@@
The training procedure of our proposed method is as shown in Algorithm 1. It begins by sampling the co-player $\pi^{\prime}$ agent that coordinates with the ego-agent from the co-player population during the training. This sampling is based on coordination performance (return), as described in Section 3.2. In the first epoch, the ego-agent coordinates with itself in the self-play method without sampling the co-player. After co-player sampling, our method employs a Replay-decision based on a Bernoulli distribution. This decision process determines whether to sample an environment $\theta$ from the co-player’s environment buffer or an unseen random environment $\theta$ from the finite set of training environments. When an environment is sampled from the co-player’s environment buffer, our method follows the replay distribution given by Equation (1), which is based on the return. It then collects the trajectories from the ego-agent and the sampled environment/co-player pairs. The ego-agent’s policy is updated whenever an environment is sampled from the co-player’s environment buffer. Coordination scores are calculated with collected trajectories and updated to the individual environment buffer of the sampled co-player accordingly. After N episodes training, the ego agent’s frozen weight policy is added to the co-player population $\mathfrak{B}$ with the environment buffer.
@@
The replay distribution is a combination of two distributions: ($P_{S}$), based on the learning potential score, which is measured using the return. ($P_{C}$), based on how long each environment was sampled from the environment buffer, as detailed in the previous study Jiang et al. (2021b). The staleness coefficient $\rho$in [0, 1] For more details on $P_{S}$, see Section 3.3.
@@
Figure 2: Evaluation Layouts. The evaluation layouts were selected based on a previous study Yang et al. (2022). These layouts are selected to cover a range of difficulties, from easy ($test_{0}$) to difficult ($test_{4}$). The selection of layouts was based on rewards, to gain insight into the human-AI coordination performance across different difficulty levels. Easier layouts allow agents to work independently within separate areas, whereas harder layouts require coordination with other agents without collisions.
@@ 3.2 Prioritized Co-player Sampling
The multiagent UED study for two-player zero-sum games Samvelyan et al. (2023) prioritized sampling the co-player with the highest-regret environment in their buffer and focused on minimizing regret to obtain a minimax regret policy. However, it is hard to apply directly to the Human-AI coordination (common pay-off game) setting. The ego-agent needs to learn a policy that not only minimizes regret but also considers the reward from coordination interaction, since the ego and the co-player agent receive the common reward together. We need to set the co-player sampling metric by considering the coordination for Human-AI coordination. Most Human-AI coordination studies Zhao et al. (2023); Lou et al. (2023) typically use return as a co-player sampling metric. They sample the co-players who have a low return on a task, which means are hard to coordinate with. Because it optimizes the lower bound of coordination performance with any co-player in the population, ensuring that the ego-agent can coordinate well with various co-players and human players Zhao et al. (2023). We propose methods that prioritize the sampling of the co-player with the lowest-return environment in their buffer. It is different from previous Human-AI coordination co-player sampling by considering the environment jointly. Each co-player environment buffer has a fixed k-size and it is stored based on the lowest return environments.
[…]
Here, $\mathfrak{B}$ is the co-player population, $\bm{\Lambda}(\pi^{\prime})$ is the environment buffer of co-player agent $\pi^{\prime}$, and $Return(\theta,\pi^{\prime})$ is the return of the environment/co-player pairs $(\theta,\pi^{\prime})$.

### v2

@@ Abstract
Zero-shot human-AI coordination is the training of an ego-agent to coordinate with humans without human data. Most studies on zero-shot human-AI coordination have focused on enhancing the ego-agent’s coordination ability in a given environment without considering the issue of generalization to unseen environments. Real-world applications of zero-shot human-AI coordination should consider unpredictable environmental changes and the varying coordination ability of co-players depending on the environment. Previously, the multi-agent UED (Unsupervised Environment Design) approach has investigated these challenges by jointly considering environmental changes and co-player policy in competitive two-player AI-AI scenarios. In this paper, our study extends a multi-agent UED approach to zero-shot human-AI coordination. We propose a utility function and co-player sampling for a zero-shot human-AI coordination setting that helps train the ego-agent to coordinate with humans more effectively than a previous multi-agent UED approach. The zero-shot human-AI coordination performance was evaluated in the Overcooked-AI environment, using human proxy agents and real humans. Our method outperforms other baseline models and achieves high performance in human-AI coordination tasks in unseen environments. The source code is available at https://github.com/Uwonsang/ACD_Human-AI
@@ 1 Introduction
Deep reinforcement learning has achieved remarkable success in diverse domains such as gaming Silver et al. (2017); Berner et al. (2019); Vinyals et al. (2019), autonomous vehicles Kiran et al. (2021), and robotic controls Fang et al. (2019). Research in deep reinforcement learning has expanded to multi-agent reinforcement learning, where multiple agents cooperate or compete to achieve specific objectives Zhang et al. (2021); Oroojlooy and Hajinezhad (2023). Most studies in this field have focused on developing powerful AI beyond human performance through AI-AI interactions. Recently, interest has shifted to problem-solving through human-AI interaction.
@@
Some human-AI interaction studies focus on training agents that coordinate with humans in common tasks and problem-solving. These studies typically require a large amount of human data for training, but acquiring sufficient human data is challenging in real-world applications. To overcome this limitation, a zero-shot human-AI coordination study Strouse et al. (2021) suggests an approach that trains the ego-agent to coordinate with humans without human data. This approach mainly uses population-based training (PBT) Strouse et al. (2021); Lupu et al. (2021); Zhao et al. (2023); Lou et al. (2023). During training, the ego-agent coordinates with various co-player agents from the population pool, rather than with a specific co-player agent. This prevents overfitting to a specific coordination strategy and helps generalize across diverse human coordination strategies.
@@
Unsupervised environment design (UED) Dennis et al. (2020) is a method that selectively curates environments for the agent by using a curriculum based on a given utility function. It curates increasingly difficult environments based on the current agent’s policy. This approach allows the agent to develop the ability to generalize across all possible environments by experiencing diverse and challenging scenarios. These methods Jiang et al. (2021a, b); Parker-Holder et al. (2022) mainly use regret as a utility function. Recently, UED methods have been extended to competitive multi-agent settings Samvelyan et al. (2023). In a multi-agent setting, the co-player policy varies depending on the environment; therefore, the multi-agent UED curates joint environment/co-player pairs for each agent. This approach has successfully trained robust agents across diverse environments and co-players. By extending this method to the zero-shot human-AI coordination setting, we could train a robust agent that can coordinate with real humans across diverse environments.
However, existing approaches in zero-shot human-AI coordination only focus on unseen co-players and do not consider generalization to unseen environments. Real-world applications of human-AI coordination face various unpredictable variables and situations, including diverse partners and changing environmental conditions. Therefore, it is essential to train the ego-agent to be robust against both diverse co-players and environments. While UED methods could potentially address this challenge by extending to zero-shot human-AI coordination settings, the existing multi-agent UED method is not directly applicable to coordination tasks. This method was developed for competitive zero-sum games and does not match the collaborative nature of human-AI coordination tasks. This mismatch motivates the need for a specialized approach that combines the benefits of zero-shot coordination and unsupervised environment design while addressing their respective limitations.
@@
Figure 1: Overview of the proposed method. The ego-agent is trained using a population of co-players whose individual environment buffers maintain a fixed size of K, containing the environments with the lowest coordination scores. Our approach samples the co-player based on their coordination scores within the population. Depending on the replay decision, the environment is either sampled from the unseen environment buffer or the co-player’s buffer, guided by a replay distribution that prioritizes environment sampling based on coordination scores. After sampling, our method calculates the return from the episode trajectory within the sampled environment, which is used to compute the coordination score. This coordination score is then updated within the co-player’s buffer.
@@
To address these problems, we introduce the automatic curriculum design for zero-shot human-AI coordination. We adopt a return-based utility function and co-player sampling for zero-shot human-AI coordination. The return measures the ego-agent’s coordination performance with co-players. It helps the agent learn better coordination strategies even in challenging scenarios and with difficult co-players. The contributions of our study are as follows: (1) We propose a multi-agent UED framework for zero-shot human-AI coordination that uses a return-based utility function and co-player sampling strategy. (2) We show that our method demonstrates higher coordination performance with human proxy agents and real human partners compared to other baselines. Figure 1 shows an overview of our proposed method.
The rest of the paper is organized as follows. Section 2 provides a review of related work in unsupervised environment design and zero-shot human-AI coordination. Section 3 presents our proposed automatic curriculum design method, including the return-based utility function and the co-player sampling strategy. Section 4 describes the experimental setup and evaluation metrics. Section 5 presents the experimental results with human proxy agents and real human participants. Finally, Section 6 concludes the paper and discusses future research directions.
@@ 2.1 Unsupervised Environment Design Strategies
Unsupervised Environment Design (UED) is a method in which the teacher curates a curriculum to enable the student agent to generalize across all possible levels of the environment based on a utility function. The simplest version of the UED teacher uses the utility function as a constant $C$ for uniform random sampling of the environment level $U_{t}(\pi,\theta)=C$. Here, $t$ is the teacher, $\pi$ is the student’s policy, $\theta$ is the environment level parameter Jakobi (1997). Most UED studies use regret as a teacher’s utility function, defined as the difference between the expected return of the current policy and the optimal student policy. The teacher aims to maximize this regret Dennis et al. (2020); Gur et al. (2021), which can be defined as : $U_{t}(\pi,\theta)=\max_{\pi^{*}\in\Pi}\{\textsc{Regret}^{\theta}(\pi,\pi^{*})\}=\max_{\pi^{*}\in\Pi}\{V_{\theta}(\pi^{*})-V_{\theta}(\pi)\}$, where $\pi^{*}$ is the optimal policy in $\theta$, $\Pi$ is the set of student policies. This regret-based utility function encourages the teacher to curate challenging environments for the student agent. Through the training process, the student’s policy will converge to the minimax regret policy by minimizing regret in teacher-curated environments Dennis et al. (2020): $\pi\in\arg\min_{\pi\in\Pi}\{\max_{\theta,\pi^{*}\in\Theta,\Pi}\{\textsc{Regret}^{\theta}(\pi,\pi^{*})\}\}$ where $\Theta$ is the set of environment- level parameters. Since the optimal policy $\pi^{*}$ cannot be accessed in practice, the regret must be approximated.
[…]
MAESTRO Samvelyan et al. (2023) extends single-agent UED to multi-agent UED in a competitive setting by curating environment/co-player pairs while considering the environment-dependent co-player policy. Recently, the Overcooked Generalisation Challenge (OGC) Ruhdorfer et al. (2024) highlighted generalization challenges in human-AI coordination, showing that existing UED-based methods struggle to achieve zero-shot coordination with novel partners and environments.
However, most existing methods are primarily effective in single-agent and competitive settings. Regret-based utility functions are not suitable for human-AI coordination tasks, where coordination and common rewards are essential, as they do not directly reflect collaborative success with diverse partners. In this paper, our method extends competitive multi-agent UED to a zero-shot human-AI coordination setting by using an appropriate utility function.
@@ 2.2 Zero-Shot Human-AI Coordination
In multi-agent cooperation tasks, achieving generalization to partner agents is an important problem. Zero-shot coordination Hu et al. (2020) and ad hoc team cooperation Stone et al. (2010) aim to address this problem. These approaches commonly rely on self-play methods in which an agent learns through collaboration with its clone. However, self-play methods suffer from limitations. Because they assume partner agents to be optimal or similar to themselves during training, their performance is poorly coordinated with unseen partners such as humans or other agents Carroll et al. (2019).
To address these issues, Devlin et al. (2011) increases partner diversity and encourages self-play agents to learn different strategies. One of these approaches is Population-Based Training (PBT), which leverages diversity within a population of partner agents. In each iteration, agents are paired with different partners in the population, allowing them to learn various strategies simultaneously. Fictitious Co-Play (FCP) Strouse et al. (2021) is a well-known PBT-based approach. It uses the population of past checkpoint self-play agents as partners to train the ego-agent. Trajectory diversity (TrajeDi) Lupu et al. (2021) improved performance in cooperative tasks by generating diverse policies within the population. It used the Jensen-Shannon divergence to measure the diversity between different policies. Similarly, MACOP Yuan et al. (2023) focused on expanding partner diversity by continually generating a wide range of incompatible teammates through evolutionary processes. Furthermore, MEP Zhao et al. (2023) was proposed to mitigate potential issues arising from changes in the distribution of self-play agents during interactions with unseen partners. Recent approaches have further advanced zero-shot coordination. E3T Yan et al. (2023) introduces an efficient end-to-end training framework using a mixture partner policy to eliminate the need for pre-trained populations. CoMeDi Sarkar et al. (2023) leverages cross-play to encourage greater diversity than traditional statistical approaches and achieve generalized cooperation.
In addition to these diversity-driven approaches, recent work has explored the use of large language models (LLMs) for human-AI coordination. HAPLAN Guan et al. (2023) improves coordination by establishing preparatory conventions using LLM before interaction. HLA Liu et al. (2024) employs a hierarchical framework that combines high-level reasoning with low-level control for real-time coordination.
However, most of these approaches focus on improving coordination with unseen partners through partner diversity and do not explicitly address the challenges posed by unseen environments. In this study, we extend the focus to coordination across diverse unseen environments by constructing a population of co-player agents trained in various settings. Through this approach, we seek to address the poor coordination between AI agents and human partners in unseen environments.
@@ 3.1 Overview
2:  for NN episodes do
3:   $\pi^{\prime}\sim\mathfrak{B}$ $\triangleright$ Sample co-player via Eq.2
@@
6:    $\theta\sim\bm{\Lambda}(\pi^{\prime})$ $\triangleright$ Sample a replay env with dist.1
@@
10:    $\theta\sim\Lambda_{unseen/train}$ $\triangleright$ Sample unseen random env
@@
In this section, we describe our proposed method, Automatic Curriculum Design for Zero-Shot Human-AI Coordination, which is an auto-curriculum that trains the ego-agent using return-based utility functions. Our method focuses on training the ego-agent to coordinate well with human partners in diverse environments. It is based on the MAESTRO algorithm Samvelyan et al. (2023), which was developed for multi-agent UED. Our method extends MAESTRO to human-AI coordination settings. It curates environment/co-player pairs with learning potential for the ego-agent in human-AI coordination.
The training procedure of our proposed method is shown in Algorithm 1. It begins by sampling the co-player agent $\pi^{\prime}$ that coordinates with the ego-agent from the co-player population during training. This sampling is based on coordination performance (return), as described in Section 3.2. In the first epoch, the ego-agent coordinates with itself using the self-play method without sampling the co-player.
After co-player sampling, our method employs a Replay-decision based on a Bernoulli distribution. This decision process determines whether to sample an environment $\theta$ from the co-player’s environment buffer or randomly sample an unseen environment $\theta$ from the finite set of training environments not yet stored in the co-player’s environment buffer. When an environment is sampled from the co-player’s environment buffer, our method follows the replay distribution given by Equation (1), which is based on the return. It then collects trajectories from the ego-agent in the sampled environment/co-player pairs. The ego-agent policy is updated whenever an environment is sampled from the co-player’s environment buffer. Coordination scores are calculated using the collected trajectories and updated in the individual environment buffer of the sampled co-player accordingly. After N episodes of training, the ego agent’s frozen weight policy is added to the co-player population $\mathfrak{B}$ with the environment buffer.
@@
The replay distribution is a combination of two distributions: ($P_{S}$), based on the learning potential score, which is measured using the return. ($P_{C}$), based on how long each environment was sampled from the environment buffer. This equation follows the previous replay mechanism PLR Jiang et al. (2021b). The staleness coefficient $\rho\in[0,1]$ is a weighting hyperparameter that balances ($P_{S}$) and ($P_{C}$). For more details on $P_{S}$, see Section 3.3.
@@
Figure 2: Evaluation Layouts. The evaluation layouts were selected based on a previous study Yang et al. (2022). These layouts are selected to cover a range of difficulties, from easy ($test_{0}$) to difficult ($test_{4}$). Easier layouts allow agents to work independently within separate areas, whereas harder layouts require coordination with other agents without collisions.
@@ 3.2 Prioritized Co-player Sampling
The multi-agent UED study for two-player zero-sum games Samvelyan et al. (2023) prioritizes sampling the co-player with the highest regret environment in its buffer and focuses on minimizing regret to obtain a minimax regret policy.
However, this approach is hard to apply directly to the human-AI coordination setting (common pay-off game). The ego-agent needs to learn a policy that not only minimizes regret but also considers the reward from coordination interaction, since the ego and co-player agent receive the common reward together. We need to set the co-player sampling metric by considering coordination for human-AI coordination. Most human-AI coordination studies Zhao et al. (2023); Lou et al. (2023) use return as a co-player sampling metric. They sample co-players who achieve low returns when coordinating with the ego-agent, which indicates challenging coordination scenarios. This approach optimizes the lower bound of coordination performance with any co-player in the population, ensuring that the ego-agent can coordinate well with various co-players and human players.
Unlike previous human-AI coordination approaches that sample co-players based on low returns in a static environment, our method jointly considers both the co-player and the environments they have encountered during training. Equation (2) formalizes this strategy by sampling the co-player whose buffer contains the lowest-return environment in the co-player population. Each co-player environment buffer has a fixed size of k and stores environments based on the lowest return. It combines two existing ideas: (1) sampling co-players with low returns to improve worst-case coordination Zhao et al. (2023); Lou et al. (2023), and (2) prioritizing high-potential learning experiences from environment buffers Jiang et al. (2021b).
$\textsc{Co-Player}\in\arg\min_{\pi^{\prime}\in\mathfrak{B}}\left\{\min_{\theta\in\bm{\Lambda}(\pi^{\prime})}Return(\theta,\pi^{\prime})\right\}$ (2)
Here, $\mathfrak{B}$ is the co-player population, consisting of ego-agent policies that are periodically frozen during training, $\bm{\Lambda}(\pi^{\prime})$ is the environment buffer of the co-player agent $\pi^{\prime}$, and $Return(\theta,\pi^{\prime})$ is the return of the environment/co-player pairs $(\theta,\pi^{\prime})$.